Given P(E) = .5 P(F) = .5 P(E AND F) = .25

Are E and F Independent?

Are E and F Mutually Exclusive?
Answer #1

P(E AND F) = 0.25
P(E) * P(F) = 0.5*0.5 = 0.25

Since P(E AND F) = P(E) *P(F) , E AND F are independent

Here, P(E and F) is not equal to 0

So, not mutually exclusive